3 gifts wrappers can wrap a box full of gifts in 20 minutes, 30 minutes and 80 minutes respectively how long will it take them to wrap the box full of gifts if they work together
Answer by rothauserc(4718)

let t be the time to complete the job, then
:
(1/20) + (1/30) + (1/80) = 1/t
:
(12/240) + (8/240) + (3/240) = 1/t
:
23/240 = 1/t
:
******************************************
t = 240 / 23 = 10.4348 approx 10.4 minutes
******************************************
: